An open cylindrical tank 4 ft in diameter and 6 ft deep is filled with water and rotated about it's axis at 60 rpm. How much liquid is spilled, and how deep is the water at the axis?

1
Expert's answer
2022-03-23T06:29:03-0400

Volume of Tank = πR²h


R = Diameter /2 = 4/2 = 2ft


h = 6ft


= πR²h


= π"\times""\times"6


= 24π ft³


Let say it can be rotated its axis at an angle 60 rpm

Then Volume of water which will not spill = πR²hCos60


= 2πCos60


=-5.98